OpenStudy (anonymous):

1. Find the 42nd term in the sequence. a1=10; d = 6 2. Find the 27th term in the sequence. 59, 56, 53, 50, ... 3. A swim pass costs $30 for the first month. Each month after that, the cost is $20 per month. Riley wants to swim for 12 months. The sequence for this situation is arithmetic. What is the first term of this sequence? 4. The following sequence is an arithmetic sequence. 10, 22, 34, 46

OpenStudy (anonymous):

L=10(42-1)6 L=10*41*6 L=2,460 ANS. FOR THE 42 nd. TERM.

OpenStudy (anonymous):

2.47, 44, 41, 38, 35, 32, 29....

OpenStudy (anonymous):

thanks :) and ok one question per post